The leftlist package^{∗}

Andrew D. Lewis

1998/04/03

1 The leftlist environment

Sometimes I want to make a single layered, numbered list but I don’t want it in-

leftlist dented like enumerate. For this I use this package with its leftlist environment. Use it like any list environment. But note it has only one level. The text in the list is indented by \parindent. So this list won’t look so good if you like to have

\parindent as zero. . .

The option alph may be used to specify a list which is numbered (a), (b),… if you prefer this to 1, 2,… The default option is arabic.

The option color may be used to make the numbers color. This is intended to be used when the command \enumcolor is defined.

2 Implementation

1 ⟨∗package⟩

‘alph’ option.

2

 3 \def\@ADL@llistcolor{\z@}
 4 \DeclareOption{alph}{\def\@ADL@alphleft{\@ne}}
 5 \DeclareOption{arabic}{\def\@ADL@alphleft{\z@}}
 6 \DeclareOption{color}{\def\@ADL@llistcolor{\@ne}}
 7 \ExecuteOptions{arabic}
 8 \ProcessOptions

9

This is straightforward.

10

11 \newcounter{leftlist}
12 \newenvironment{leftlist}{\list{\labelleftlist}{\itemindent\z@
13 \ifnum\@ADL@llistcolor=\@ne
14  \renewcommand{\makelabel}[1]{{\enumcolor##1}\hfil}
15 \else
16  \renewcommand{\makelabel}[1]{##1\hfil}

∗This file has version number v1.0b, last revised 1998/04/03.

17 \fi
18 \labelwidth\parindent
19 \advance\labelwidth by-\labelsep
20 \leftmargin\labelwidth\advance\leftmargin\labelsep
21 \topsep 3\p@ plus \p@ minus \p@\parsep 1.5\p@ plus \p@ minus .5\p@
22 \itemsep\parsep\partopsep \p@ plus \p@ minus \p@
23 \usecounter{leftlist}}}
24 {\endlist}
25 \ifnum\@ADL@alphleft=\@ne
26 \renewcommand\theleftlist{\alph{leftlist}}
27 \newcommand\labelleftlist{(\theleftlist)}
28 \else
29 \renewcommand\theleftlist{\arabic{leftlist}}
30 \newcommand\labelleftlist{\theleftlist.}
31 \fi

32

33 \newenvironment{alphalist}{\@enumdepth=1\enumerate}{\endenumerate}

34

35 ⟨/package⟩